Application of Overhead; Schedule of Cost of Goods Manufactured Haughton Company uses a job costing system for its production costs and a predetermined factory overhead rate based on direct labor costs to apply factory overhead to all jobs. During the month of July, the firm processed three jobs: X13, X14, and X15, of which X13 was started in June. Inventories July 1 July 31 Direct materials Work in process Finished goods 36,500 41,000 Cost of Goods Sold, July Direct materials purchased in July 55,000 Issued to production: X13 X14 X15 16,380 24,220 14,000 Factory labor-hours used ($30/hour X13 X14 X15 Indirect labor 3,500 2,800 1,600 6,900 Other factory overhead costs incurred: Rent Utility Repairs and maintenance Depreciation Other $131,500 180,600 188,500 31,100 56,000 As of July 31, Job X13 was sold and Jobs X14 and X15 were still in processing. Total factory overhead applied in July was $900,600. Required 1. Compute the predetermined factory overhead rate. 2. Compute the amount of materials inventory at the end of July. 3. Compute the actual factory overhead cost incurred during the month of July 4. Compute the ending balance of the work-in-process inventory account for July. 5. Prepare the statement cost of goods manufactured for July 6. Compute the amount of over- or underapplied overhead 7. What is the cost per unit of Job X13 if it has a total of 100 units? 8. Prepare the statement of cost of goods sold for July
Answer #1

1)Actual Direct labor hours worked during July = 3500 +2800 +1600 = 7900

Actual Direct labor cost = 7900 *30 = 237000

Predetermined overhead rate= Overhead applied in July /Actual direct labor cost

              = 900600/ 237000

                 = 3.8 or 380% of direct labor cost

2)Amount of material inventory at end of July =Beginning material+ Purchase- issued for production

           = 36500+ 55000- [16380+24220+14000]

           =36500+ 55000 - 54600

          = 36900

3)Actual Factory overhead Incurred = Indirect Labor + other overhead

                 = 6900+131500+180600+188500+131100+56000

                 = 694600

4) Cost of Job X14 : 24220 Direct material + (2800*30 = 84000)Direct labor + (84000*380%) overhead applied

            = 427420

Cost of Job X 15 : 14000 + (1600*30= 48000)+(48000*380%)

       = 14000+ 48000+182400

= 244400

WIP at end of July = 427420 +244400

        = 671820